OpenStudy (anonymous):

Given that "a" is an odd multiple of 7767, find the greatest common divisor of 6a^2+49a+108 and 2a+9.

OpenStudy (anonymous):

how is that?... can you detail?

OpenStudy (queelius):

a=(2k+1)*7767, so we want the gcd of: 6[(2k+1)*7767]^2+49(2k+1)7767 = (2k+1)*7767*(6*7767*(2k+1) + 49) and 2(2k+1)7767 Notice that they both have (2k+1)*7767 in common, and that (6*7767*(2k+1) + 49) is an odd number. Thus, the gcd is (2k+1)*7767.

OpenStudy (queelius):

Ooops, I didn't write down the constants in both equations. So, I solved the wrong problem.

OpenStudy (queelius):

Let's try again, maybe this time with your help.

OpenStudy (anonymous):

sure.

OpenStudy (queelius):

Those constants throw a monkey wrench into the works! :)

OpenStudy (anonymous):

the answer given by loser66 is 9... but I'm waiting for explanation. not sure...if 9 is the greatest GCD.

OpenStudy (anonymous):

Ok, a can be 9(the greatest) but how do you find it?

OpenStudy (queelius):

Ok, I found it.

OpenStudy (queelius):

So, first, expand the two equations.

OpenStudy (queelius):

We are given that equation one is: 6[(2k+1)7767]^2+49(2k+1)*7767+108

OpenStudy (queelius):

and equation two is: 2*(2k+1)*7767+9

OpenStudy (loser66):

ok , we can do together

OpenStudy (loser66):

6a^2 +49a+108 = (2a+9)(3a+11)+9

OpenStudy (queelius):

If we multiply these out, we get: 1447830936 k^2+1448592102 k+362338425 and 31068 k+15543

OpenStudy (loser66):

Now, we consider 2a+9 = ?? 9

OpenStudy (queelius):

Applying the gcd algorithm to all of these numbers gets you a gcd of 9.

OpenStudy (loser66):

Not that

OpenStudy (queelius):

That means you can "pull" out that 9 in both, and thus their gcd is 9.

OpenStudy (loser66):

2a+9 =?? 9 if a = 7761, we have 9| 2a+9,

OpenStudy (loser66):

so, no matter how a factor of a is, 9|2a+9

OpenStudy (loser66):

I know, a is an odd multiple of 7767, so the bottom line is a = 7767 and this guy divided by 9

OpenStudy (queelius):

We get the same answer, but I'm not following how you arrived at yours. For instance, I don't know what 9|2a+9 means.

OpenStudy (queelius):

I didn't get 2a+9 as the gcd, I got 9. What was the expected answer danny?

OpenStudy (queelius):

But I think I like your approach better, so you may be right.

OpenStudy (loser66):

nope, don't trust me, I don't know number theory!!

OpenStudy (anonymous):

Oh! I got it! We can use the Euclidean Algorithm. gcd(6a^2+49a+108,2a+9) =gcd(6a^2+49a+108-(2a+9)(3a+11),2a+9) =gcd(6a^2+49a+108-(6a^2+49a+99),2a+9) =gcd(9,2a+9). Since both 2a and 9 are multiples of 9, 2a+9 is also a multiple of 9 so the greatest common divisor is 9.

OpenStudy (queelius):

Cool.

OpenStudy (queelius):

Good work.

OpenStudy (loser66):

hehehe... the Asker is better than the helpers!!! how ironic situation is!!

OpenStudy (anonymous):

no, that's not the way, your answer helped me.